Mark put $2748.00 in his savings account. The bank he uses calculated his interest using a simple interest formula. Mark left the money in the bank for 2 years. At the end of the 2 years, he had $2967.84 in his savings account. What was the interest rate the bank used?
Answer by sarah_adam(201)

The formula for finding simple interest is: Interest = Principal * Rate * Time
Principal = $2748
Time = 2years
R = Rate of interest
Interest = Money at the end of two years - Initial amount
Interest = 2967.84 - 2748 = $219.84
219.84 = 2748 * R * 2
R = (219.84)/2*2748 = 0.04
Therfore the interest bank gave was 4% (0.04*100 = 4)